True or False? A clinical trial is conducted to compare an experimental medication to placebo to reduce the symptoms of asthma. Two hundred participants are enrolled in the study and randomized to receive either the experimental medication or placebo. The primary outcome is self-reported reduction of symptoms. Among 50 participants who receive the experimental medication, 30 report a reduction of symptoms as compared to 15 participants of 50 assigned to the placebo. The 95% CI for the relative risk of participants reporting a reduction of symptoms between the experimental and placebo groups is between 1.529 and 8.012.
Answer #1

The table showing the numbers based on the study is as follows:

Reduction No Reduction Total
Medication 30 20 50
Placebo 15 35 50

RR: Relative Risk = (30/50) / (15/50) = 2 (Assuming the reduction of of symptoms as incidence in each of the medication and placebo groups)

The 95% confidence interval for ln(RR) is given by - :

ln(2) +-   1.96(sqrt [ ((20/30)/50) + ((35/15)/50) ]

=ln(2) +- 0.48

= 0.693    +- 0.480

= (0.693-0.480 , 0.693+0.480)

=(0.213, 1.173)

Thus the confidence interval for RR is

antilog of the above values = (1.237, 3.232)

Thus the statement mentioned is false as the values don't match